# Efficient Compositional Multi-tasking for On-device Large Language Models

### v1

@@ Appendix D Qualitative Analysis
In contrast, Learnable Calibration++ demonstrates significantly better performance, successfully executing both tasks in the compositional settings.
We conducted further qualitative analysis by examining outputs from different methods for the same input. This analysis focused on the cross-lingual summarization task, with results shown in Figures A7, A8, A9, and A10.
@@
Cross-Lingual (English to Spanish) Summarization Reference: Eric y Rob van a ver un stand-up en YouTube. Translated reference: Eric and Rob are going to watch a stand-up on YouTube. Zero-shot prediction: Eric: Máquina! Zero-shot translated prediction: Eric: Machine! Main-task LoRA prediction: Eric: Máquina! Main-task LoRA translated prediction: Eric: Machine! Auxiliary-task LoRA prediction: Eric: Máquina! Rob: ¡Esto es genial! Eric: Lo sé. Y muestra cómo los estadounidenses ven a Rusia. Rob: Y es muy divertido. Eric: Lo sé. Y me encanta la parte del tren. Rob: ¡Hahahaha! No se habla a la máquina así. Eric: ¿Y es su única actuación? Rob: No sé. Lo veré. Eric: Sí. Rob: ¡Sí! Auxiliary-task LoRA translated prediction: Eric: Machine! Rob: This is great! Eric: I know. And it shows how Americans see Russia. Rob: And it’s really fun. Eric: I know. And I love the part of the train. Rob: Hahaha! You don’t talk to the machine like that. Eric: And it’s your only performance? Rob: I don’t know. I’ll see it. Eric: Yes. Rob: Yes! Linear merge prediction: Eric: Máquina! Linear merge translated prediction: Eric: Machine! Multi-step LoRA usage prediction: Eric y Rob hablan de un video divertido que muestra una máquina hablando como estadounidenses. Lo encuentran divertido y planean ver más de los videos de stand-up de Eric. Multi-step LoRA usage translated prediction: Eric and Rob talk about a funny video that shows a machine talking like Americans. They find it fun and plan to see more of Eric’s stand-up videos. Joint-expert LoRA prediction: Eric y Rob están viendo un stand-up de Eric, un hombre americano que hace stand-up. Eric está muy divertido y se ríe mucho. Rob también está viendo y se ríe. Joint-expert LoRA translated prediction: Eric and Rob are watching a stand-up by Eric, an American man who makes stand-up. Eric is very funny and laughs a lot. Rob is also watching and laughing. Learnable Calibration prediction: Rob y Eric están interesados en ver los vídeos de stand-up de Eric. Learnable Calibration translated prediction: Rob and Eric are interested in watching Eric’s stand-up videos. Learnable Calibration++ prediction: Eric y Rob están viendo los stand-up de Eric. Learnable Calibration++ translated prediction: Eric and Rob are watching Eric’s stand-up.
@@
Appendix E Learning Curves Analysis of Our Approach

### v2

@@ Appendix G Qualitative Analysis
In contrast, Learnable Calibration++ demonstrates significantly better performance, successfully executing both tasks in the compositional task.
We conducted further qualitative analysis by examining outputs from different methods for the same input. This analysis focused on the cross-lingual summarization task, with results shown in Figures A8, A9, A10, and A11.
@@
Cross-Lingual (English to Spanish) Summarization Reference: Eric y Rob van a ver un stand-up en YouTube. Translated reference: Eric and Rob are going to watch a stand-up on YouTube. Zero-shot prediction: Eric: Máquina! Zero-shot translated prediction: Eric: Machine! Main-task LoRA prediction: Eric: Máquina! Main-task LoRA translated prediction: Eric: Machine! Auxiliary-task LoRA prediction: Eric: Máquina! Rob: ¡Esto es genial! Eric: Lo sé. Y muestra cómo los estadounidenses ven a Rusia. Rob: Y es muy divertido. Eric: Lo sé. Y me encanta la parte del tren. Rob: ¡Hahahaha! No se habla a la máquina así. Eric: ¿Y es su única actuación? Rob: No sé. Lo veré. Eric: Sí. Rob: ¡Sí! Auxiliary-task LoRA translated prediction: Eric: Machine! Rob: This is great! Eric: I know. And it shows how Americans see Russia. Rob: And it’s really fun. Eric: I know. And I love the part of the train. Rob: Hahaha! You don’t talk to the machine like that. Eric: And it’s your only performance? Rob: I don’t know. I’$l$ see it. Eric: Yes. Rob: Yes! Linear merge prediction: Eric: Máquina! Linear merge translated prediction: Eric: Machine! Multi-step LoRA usage prediction: Eric y Rob hablan de un video divertido que muestra una máquina hablando como estadounidenses. Lo encuentran divertido y planean ver más de los videos de stand-up de Eric. Multi-step LoRA usage translated prediction: Eric and Rob talk about a funny video that shows a machine talking like Americans. They find it fun and plan to see more of Eric’s stand-up videos. Joint-expert LoRA prediction: Eric y Rob están viendo un stand-up de Eric, un hombre americano que hace stand-up. Eric está muy divertido y se ríe mucho. Rob también está viendo y se ríe. Joint-expert LoRA translated prediction: Eric and Rob are watching a stand-up by Eric, an American man who makes stand-up. Eric is very funny and laughs a lot. Rob is also watching and laughing. Learnable Calibration prediction: Rob y Eric están interesados en ver los vídeos de stand-up de Eric. Learnable Calibration translated prediction: Rob and Eric are interested in watching Eric’s stand-up videos. Learnable Calibration++ prediction: Eric y Rob están viendo los stand-up de Eric. Learnable Calibration++ translated prediction: Eric and Rob are watching Eric’s stand-up.
@@
Appendix H Learning Curves Analysis of Our Approach